1000 people attended the Everett Aquasox Minor League Fan Appreciation baseball game. The total receipts for that game were $2670. Adult ticket prices were $3 and children ticket prices were S1.50. How many adults and how many children attended this game?
Answer #1

Let

number of adult=x

hence number of children = 1000 -x

Since total revenue is 2670, set up as

3x +1.50(1000-x) = 2670

3x+1500−1.50x=2670

1.5x+1500=2670

1.5x=2670−1500

1.5x=1170

x=1170/1.5

x = 780

Thus,

Adult = 780

children = 1000 -780 =220